OpenStudy (anonymous):

x^2-3x=-5/4 Which choices are solutions to the following equation?

OpenStudy (anonymous):

3/2 (3/2)^2=9/4 (x-3/2)^2=(-5/4)+(9/4) (x-3/2)^2=-1 (x-3/2)=(-1)^2 x=(3/2)±1 x=5/2 or x=1/2